1. An alpha particle is sent directly toward a target nucleus containing 92 protons. The alpha 

particle has an initial kinetic energy of 0.48pJ. What is the least center to center distance 

the alpha particle will from the target nucleus (assuming the nucleus does not move)?

2. Three +0.12C charges from an equilateral triangle 1.7m on a side. Using energy supplied

At the rate of 0.83KW, how many days would be required to move one of the charges to 

the midpoint of the line joining the other two charges?


3. The electric potential difference between the ground and a cloud in a particular 

thunderstorm is 1.2X109 V. In the unit of electron- volt, what is the magnitude of the 

electric potential energy of an electron that moves between ground and the clouds?




1
Expert's answer
2021-11-11T17:07:34-0500

1)

"E=\frac{kq_1q_2}r,\implies"

"r=\frac{kq_1q_2}E=8.8\cdot 10^{-14}~m."

2)

"W=\frac{2kq^2}{tl\sqrt 3},\implies"

"t=\frac{2kq^2}{Wl\sqrt 3}=1.2^d."

3)

"\varphi=1.2\cdot 10^9\cdot 1.6\cdot 10^{-19}=1.92\cdot 10^{-10}~eV."